A 1litre vessel which is equipped with a movable piston is filled with a 2M aqueous solution of h2O2.the h2O2 decomposes to h2O and O2 is a first order process with half life 5 hours at 300K.as gas formed the piston moves upward against external pressure of 1atm.calculate work done by the gas from the start of 6hours till the end of 15 hours
1
Expert's answer
2020-11-04T14:22:20-0500

2H2O2 = 2H2O + O2

k = ln(2)/5 = 0.139 h-1

ln(C0/C) = kt

6 hours:

ln(2/C) = 0.139*6

C = 0.869

15 hours:

ln(2/C) = 0.139*15

C = 0.249

"\Delta"CH2O2 = 0.869-0.249 = 0.620 M

"\Delta"nH2O2 = 0.620 mol

"\Delta"nO2 = 0.310 mol

"\Delta"VO2 = "\Delta"nO2*R*T/p = 0.310*8.314*300/101325 = 0.00763 m3

W = p*"\Delta"VO2 = 101325*0.00763 = 773 J